OpenStudy (abdela25):

What is the f(-5) for the function f(x) = -9x-3? medals will be awarded for first best answer

OpenStudy (anonymous):

Just put -5 in for x, and you get: f(-5)=-9(-5)-3 f(-5)=45-3 f(-5)=42

OpenStudy (abdela25):

thank you very much @zeinas. medal was awarded

OpenStudy (anonymous):

You're welcome :), and thanks for the medal